A fixed asset with a cost of $30,000 and accumulated depreciation of $28,500 is sold for $3,500. What is the amount of the gain or loss on disposal of the fixed asset? Group of answer choices $2000 loss $1,500 loss $3,500 gain $2,000 loss
Answer #1

Book value of assets = Cost-Accumulated depreciation = 30000-28500 = 1500

Gain on Sale = Sale value-Book value = 3500-1500 = 2000 Gain

So answer is $2000 Gain